A hydrogen-oxygen fuel cell uses 2000dm3 of hydrogen measured at room temperature and pressure.Calculate the volume of oxygen, measured at room temperature and pressure.
1
Expert's answer
2020-07-14T15:19:17-0400

At room temperature and pressure means that;

"Temperature\ =25^0\ C\ or\ (273+25)=298K"

"Pressure\ =1\ atmospere\ of\ pressure"

Thus, the formula below is used to calculate the volume of oxygen


"\dfrac{P1V1}{T1}=\dfrac {P2V2}{T2}"


"\dfrac{1\ *\ 2000}{298}=\ \dfrac{1\ *V2}{298}"

"V2=2000"

The volume of oxygen is "2000dm^3" measured at room temperature and pressure.